# Accelerated Ostwald ripening by chemical activity

### v1

@@ Acknowledgements
We thank Job Boekhoven and Alejandro Martínez-Calvo for illuminating discussions. B.S. acknowledges support from the Princeton Center for Theoretical Science. This project has been made possible in part by grant number DAF2024-342781 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation, and NIH Grant R01GM140032.
@@
Most of the generic conclusions drawn in the main text for the equilibrium properties of phase-separating systems are motivated by the Flory-Huggins (FH) model [65, 66]. In this section, we first present the general thermodynamic principles underlying phase separation of a binary fluid (sticky and inert solvent). We proceed to connect the surface tension to the mesoscopic field theory considered in the main text. We then show concrete results for the FH free energy to provide intuition and to further support the assumptions concerning the volume fractions of each phase. We conclude by extending all these results to the case of a ternary fluid (sticky, nonsticky, and inert solvent) which is the case studied in the main text.
@@
where $\phi_{i}=N_{i}/(V\varrho)$ is the volume fraction of the $i$th species. Thus, $\phi_{\mathrm{o}}$ is uniquely determined by $\phi_{\mathrm{s}}$, as well as $N_{\mathrm{o}}$ by $N_{\mathrm{s}}$ and $V$. Furthermore, we assume that the temperature $T$ is also held constant. Upon phase separation, only $V$ and $N_{\mathrm{s}}$ vary, and so we will suppress the dependence of the free energy $F$ on the constants $T$ and $\varrho$.
@@
In the main text, the liquid phase is assumed to be finite sized and spherical with radius $R$. As a result, we must account for the surface tension $\sigma$, which leads to a curvature-induced Laplace pressure [28], $2\sigma/R$. Therefore, The coexistence volume fractions modify to $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ and $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$, which satisfy
@@
where, using Eq. (S3b), we identified the isothermal compressibility coefficient [36] $\beta(\phi_{\mathrm{s}})\equiv-(1/V)(\partial V/\partial P)_{N_{\mathrm{s}}}=1/[\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}}^{2}f^{\prime\prime}(\phi_{\mathrm{s}})]$. Indeed, the higher the surface tension $\sigma$ and compressibility $\beta$ (the susceptibility of each phase to compression due to the extra Laplace pressure), the larger is the increase in sticky volume fraction due to Laplace pressure. In the derivations to follow, only the gas-phase sticky Laplace-pressure correction will appear explicitly, which is why we write for brevity $\lambda_{\mathrm{s}\mathrm{g}}\equiv\lambda$ in the main text.
To conclude, for completeness, we present the following common simplification of the above: In a typical system, the liquid phase is much denser than the gas phase, $\phi_{\mathrm{s}\mathrm{l}}^{\circ}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq\phi_{\mathrm{s}\mathrm{l}}^{\circ}$, the former is incompressible, $\beta(\phi_{\mathrm{s}\mathrm{l}}^{\circ})\simeq 0$, and the latter is ideal, $\beta(\phi_{\mathrm{s}\mathrm{g}}^{\circ})\simeq 1/(\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}\mathrm{g}}^{\circ})$. Thus, the liquid typically has no Laplace-pressure correction, $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\simeq\phi_{\mathrm{s}\mathrm{l}}^{\circ}$ whereas the gas follows the Ostwald-Freundlich equation [67],
@@
In our coarse-grained view of a phase-separating system, a given particle configuration is modeled via a volume-fraction field. To leading order in small-wavelength excitations, we promote the homogeneous free energy (Eq. (S2)) into a free energy functional,
@@
where $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})=\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})$ is the chemical potential deep in each bulk. We multiply each side of the equation by $\mathrm{d}\phi_{\mathrm{s}}(x)/\mathrm{d}x$, and integrate from $\infty$- ∞ to $x$,
@@
where $k_{\mathrm{B}}T(-\chi)$ accounts for the attractive ($\chi>0$) energy between sticky particles. We assume that the molecular volumes of the species [1], as well as their polymeric lengths [69], are equal. The chemical potential and pressure are given by
@@
We find the volume fractions of each phase in coexistence. Phase separation occurs in this model for $\chi>2$ [69]. To model strong phase separation, $1-\phi_{\mathrm{s}\mathrm{l}}^{\circ},\phi_{\mathrm{s}\mathrm{g}}^{\circ}\ll 1$, we consider $\chi\gg 1$. Upon solving Eqs. (S4) for the FH free energy under the ansatz $\phi_{\mathrm{s}\mathrm{l}}^{\circ}\simeq 1$ and $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq 0$, we find
@@
where corrections are of order $\chi e^{-2\chi}$, which is thus consistent with $\chi\gg 1$. Inserting these volume fractions in Eq. (S7), we find
@@
with, again, corrections of order $\chi e^{-2\chi}$. Interestingly, although the FH liquid phase is much less compressible than the gas phase, $\varrho k_{\mathrm{B}}T\beta(1-e^{-\chi})=e^{-\chi}\ll\varrho k_{\mathrm{B}}T\beta(e^{-\chi})=e^{\chi}$, the Laplace-pressure corrections are equal.
@@
All four volume fractions (as well as the four Laplace-pressure corrections) can be deduced by imposing material conservation across both phases when phase separation is complete, $\zeta_{\mathrm{l}}\phi_{\mathrm{s}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{s}\mathrm{g}}^{\circ}=N_{\mathrm{s}}/(V\varrho)$ and $\zeta_{\mathrm{l}}\phi_{\mathrm{ns}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=N_{\mathrm{ns}}/(V\varrho)$, where $\zeta_{\mathrm{l}}=1-\zeta_{\mathrm{g}}$ is the volume fraction of the dense phase (total five equations with five unknowns $\{\phi_{i\alpha}^{\mathrm{L}}\}$ and $\zeta_{\mathrm{l}}$), or by flux matching during droplet growth (Eqs. (5) and (12) of the main text, for a total of five equations with five unknowns $\{\phi_{i\alpha}^{\mathrm{L}}\}$ and $\mathrm{d}R/\mathrm{d}t$).
@@
Note that Eq. (S25b) implies that $\phi_{\mathrm{ns}}(x)$ simply spatially rearranges in accord with $\phi_{\mathrm{s}}(x)$, so the positional dependence of the former is strictly acquired from the latter, $\phi_{\mathrm{ns}}(x)\to\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x))$. By multiplying each of Eqs. (S25) by $\mathrm{d}\phi_{i}(x)/\mathrm{d}x$, summing the two equations, and integrating the result from $\infty$- ∞ to $x$, we obtain
$f(\phi_{\mathrm{s}}(x),\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x)))-\frac{\kappa_{\mathrm{s}}}{2}\left[\frac{\mathrm{d}\phi_{\mathrm{s}}(x)}{\mathrm{d}x}\right]^{2}-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{s}}(x)-\phi_{\mathrm{s}\mathrm{g}}^{\circ}]-\frac{\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x))-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}],$ (S26)
Via a similar mechanism to Sec. SI.2, the surface tension can be found as
@@
Thus, per the Gibbs phase rule, only one unknown $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$(and its leading-order finite-droplet correction, $\lambda_{\mathrm{ns}\mathrm{g}}$) remain to be determined. For brevity, since we will find $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\ll 1$, we set $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\ll e^{-\chi}$ to be simply $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}=0$.
Lastly, we confirm that the FH surface tension (Eq. (S13)) does not change due to “labeling” inert (nonsticky) species. First, Eq. (S25b) simplifies for the FH free energy as $\phi_{\mathrm{ns}}(x)/[1-\phi_{\mathrm{s}}(x)-\phi_{\mathrm{ns}}(x)]=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}/(1-\phi_{\mathrm{s}\mathrm{g}}^{\circ}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ})$, which is the position-dependent analogue of the inert proteome/nonsticky protein equivalence statement, Eq. (S30). Inserting it in Eq. (S27) with the FH free energy (Eq. (S28)) results in Eq. (S13) with the binary-fluid’s FH free energy (Eq. (S14)). Thus, indeed the surface tension is ignorant of inert-protein labels, so $\sigma$ is still given by Eq. (S18).
@@
In this section, we detail the full derivation of the droplet growth laws in the presence of reactions. Particularly, we elaborate further on our approximations, list the phenomena that might occur in the liquid phase, derive the flux-matching equation (Eq. (21)) of the main text, extract $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ from it in the different limits of droplet size $R$, and from it the radius-growth laws. Throughout, we refer to the results of the FH free energy to support general intuition.
@@
Here, we construct the reaction-diffusion equations of the main text (Eq. (13)). We denote with vectors and matrices
@@
First, following linear-irreversible thermodynamics, the volume-fraction fields $\bar{\phi}(\mathbf{r},t)$ respond linearly to chemical potential gradients $\delta F/\delta[\varrho\bar{\phi}(\mathbf{r},t)]$ via a (normalized) mobility matrix denoted $\varrho k_{\mathrm{B}}T\bar{\bar{\Gamma}}(\bar{\phi}(\mathbf{r},t))$. Thus, the diffusive part of the dynamics is given by
@@
where $\kappa_{ij}=\delta_{i\mathrm{s}}\delta_{j\mathrm{s}}\kappa_{\mathrm{s}}$. For a single-component system, this reproduces Eq. (2) of the main text.
@@
Upon combining both reactions and diffusion, we find Eq. (13) of the main text,
@@
In the main text, we assume a one-to-one stoichiometric ratio between $\mathrm{s}$ and $\mathrm{ns}$ for the interconversion reactions. Here, we proceed with a general ratio.
Since $\mathrm{s}$ is created from $\mathrm{ns}$ and vice versa, $\upsilon_{\mathrm{s}}$ and $\upsilon_{\mathrm{ns}}$ must be of opposite signs. Without loss of generality (up to rescaling of $s(\bar{\phi})$), we choose $\upsilon_{\mathrm{s}}<0$ and $\upsilon_{\mathrm{ns}}=1$; $|\upsilon_{\mathrm{s}}|$ thus determines the stoichiometric ratio. (In the main text, $\upsilon_{\mathrm{s}}=-1$.) For later convenience, we also define $\bar{\upsilon}_{\perp}$ as the orthogonal vector to $\bar{\upsilon}$, so $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\upsilon}=0$. Again without loss of generality, we choose it as $\bar{\upsilon}_{\perp}=(1,|\upsilon_{\mathrm{s}}|)^{\mathrm{T}}$ (whereas $\bar{\upsilon}=(-|\upsilon_{\mathrm{s}}|,1)^{\mathrm{T}}$).
@@
SII.2 Interim simplifications
To make analytical progress from Eq. (S37), we made a number of standard simplifications in the main text [1], valid at a late stage of coarsening. We quickly recap them: First, the droplets are sparse (the interdroplet distances are much greater than their radii), so we may treat a single phase-separated droplet as if it is in isolation, diffusively exchanging material with a slightly-supersaturated dilute phase, $0<\varphi^{\infty}_{\mathrm{s}}-\phi_{\mathrm{s}}^{\circ}\ll 1$. Owing to surface tension, the droplets are spherical, with radius $R(t)$ much larger than the interface width $\xi$ (defined in Sec. SI.3). With spherical symmetry, we split the volume fraction fields to droplet interior (liquid) and exterior (gas), $\bar{\phi}_{\mathrm{l}}(r,t)=\bar{\phi}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|<R(t),t)$ and $\bar{\phi}_{\mathrm{g}}(r,t)=\bar{\phi}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|>R(t),t)$, respectively, where $\mathbf{r}_{\mathrm{CM}}$ is the droplet-center coordinate. Assuming the reactions are slow enough not to interfere with surface thermodynamics [39], the two volume fractions are always equilibrated at the interface, $\bar{\phi}_{\alpha}(R(t),t)=\bar{\phi}_{\alpha}^{\mathrm{L}}(R(t))$ ($\alpha=\mathrm{l},\mathrm{g}$). Lastly, the droplets grow quasistatically, whereby the volume fraction fields $\{\bar{\phi}_{\alpha}(r;R)\}$ are found for a quasi-steady state of a fixed radius $R$, and then droplet growth rate is inferred from the resulting steady diffusive flux.
In what follows, we present a more thorough treatment of the last approximation we have made in the main text — linearizing Eq. (S37) deep in each bulk. We now present the general scheme, and in subsequent sections we solve the resulting equations, discuss their validity in each phase, and make adjustments accordingly.
@@
For brevity, we will omit the reminder of the parametric dependence on $R$ henceforth. Note that, in accord with Sec. SI.4, the volume fractions $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ ($\simeq 1$), $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$ ($\ll 1$), and $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}$ ($\to 0$) are known explicitly, whereas $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is still unknown. As mentioned there and in Sec. III A of the main text, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ will be determined by equating the two equations for $\mathrm{d}R/\mathrm{d}t$, obtained from the sticky and nonsticky fluxes (Eqs. (5) and (12) of the main text, and Sec. SII.5 below). To obtain the equation for $\mathrm{d}R/\mathrm{d}t$, we regard $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ as a known parameter, and express the flux $J_{\mathrm{ns}}$ in term of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. The linearization of Eq. (S38) is thus an ansatz for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ — that $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\simeq\varphi_{\mathrm{ns}}^{\infty}$ — which we make to be able to compute the flux $J_{\mathrm{ns}}$ in terms of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. Once all calculations are complete and $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is found, we confirm that indeed the resulting $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is close to $\varphi_{\mathrm{ns}}^{\infty}$.

### v2

@@ Acknowledgements
We thank Job Boekhoven and Alejandro Martínez-Calvo for illuminating discussions. B.S. acknowledges support from the Princeton Center for Theoretical Science. This project has been made possible in part by the Chan Zuckerberg Initiative DAF — an advised fund of Silicon Valley Community Foundation — grant number DAF2024-342781, NIH grant number R01GM140032, and NSF grant number MCB-2410354.
@@
Most of the generic conclusions drawn in the main text for the equilibrium properties of phase-separating systems are motivated by the Flory-Huggins (FH) model [56, 57]. In this section, we first present the general thermodynamic principles underlying phase separation of a binary fluid (sticky solute and inert solvent). We proceed to connect the surface tension to the mesoscopic field theory considered in the main text. We then show concrete results for the FH free energy to provide intuition and to further support the assumptions concerning the volume fractions of each phase. We conclude by extending all these results to the case of a ternary fluid (sticky, nonsticky, and inert solvent) which is the case studied in the main text.
@@
where $\phi_{i}=N_{i}/(V\varrho)$ is the volume fraction of the $i$th species. For brevity, we refer to $\phi_{i}$ as a unitless concentration in the main text. Thus, $\phi_{\mathrm{o}}$ is uniquely determined by $\phi_{\mathrm{s}}$, as well as $N_{\mathrm{o}}$ by $N_{\mathrm{s}}$ and $V$. Furthermore, we assume that the temperature $T$ is also held constant. Upon phase separation, only $V$ and $N_{\mathrm{s}}$ vary, and so we will suppress the dependence of the free energy $F$ on the constants $T$ and $\varrho$.
@@
To treat a droplet, the liquid phase is assumed to be finite sized and spherical with radius $R$. As a result, we must account for the surface tension $\sigma$, which leads to a curvature-induced Laplace pressure, $2\sigma/R$ [26]. Thus, The coexistence volume fractions modify to $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ and $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$, satisfying
@@
where, using Eq. (S3b), we identified the isothermal compressibility coefficient [27],
$\beta(\phi_{\mathrm{s}})\equiv-\frac{1}{V}\left(\frac{\partial V}{\partial P}\right)_{N_{\mathrm{s}}}=\frac{1}{\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}}^{2}f^{\prime\prime}(\phi_{\mathrm{s}})}.$ (S8)
Indeed, the higher the surface tension $\sigma$ and compressibility $\beta$ (the susceptibility of each phase to compression due to the extra Laplace pressure), the larger is the increase in sticky volume fraction due to Laplace pressure.
For completeness, we conclude by presenting the following common simplification of Eq. (S7): In a typical system, the liquid phase is much denser than the gas phase, $\phi_{\mathrm{s}\mathrm{l}}^{\circ}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq\phi_{\mathrm{s}\mathrm{l}}^{\circ}$, and furthermore the latter can be approximated as ideal, $\beta(\phi_{\mathrm{s}\mathrm{g}}^{\circ})\simeq 1/(\varrho k_{\mathrm{B}}T\phi_{\mathrm{s}\mathrm{g}}^{\circ})$. Thus, the gas-phase volume fractions follow an Ostwald-Freundlich-type equation,
@@
In our coarse-grained view of a phase-separating system, a given particle configuration is modeled via a volume-fraction field. To leading order in small-wavelength excitations, we promote the homogeneous free energy (Eq. (S2)) into a free energy functional [58, 59],
@@
where $\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ})=\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{l}}^{\circ})$ is the chemical potential deep in each bulk. We multiply each side of the equation by $\mathrm{d}\phi_{\mathrm{s}}(x)/\mathrm{d}x$, and integrate from -$\infty$ to $x$,
@@
where $k_{\mathrm{B}}T(-\chi)$ accounts for the attractive ($\chi>0$) energy between sticky particles. We assumed for simplicity that the molecular volumes of the species [1], as well as their polymeric lengths [60], are equal. The chemical potential and pressure are given by
@@
We find the volume fractions of each phase in coexistence. Phase separation occurs in this model for $\chi>2$ [60]. To model strong phase separation, $(1-\phi_{\mathrm{s}\mathrm{l}}^{\circ}),\phi_{\mathrm{s}\mathrm{g}}^{\circ}\ll 1$, we consider $\chi\gg 1$. Upon solving Eqs. (S4) for the FH free energy under the ansatz $\phi_{\mathrm{s}\mathrm{l}}^{\circ}\simeq 1$ and $\phi_{\mathrm{s}\mathrm{g}}^{\circ}\simeq 0$, we find
@@
where corrections are of order $\chi e^{-2\chi}$, which are thus consistent with $\chi\gg 1$. Inserting these volume fractions in Eq. (S7), we find
@@
with, again, corrections of order $\chi e^{-2\chi}$.
@@
All four volume fractions (as well as the four Laplace-pressure corrections) can be deduced by imposing material conservation across both phases when phase separation is complete, $\zeta_{\mathrm{l}}\phi_{\mathrm{s}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{s}\mathrm{g}}^{\circ}=N_{\mathrm{s}}/(V\varrho)$ and $\zeta_{\mathrm{l}}\phi_{\mathrm{ns}\mathrm{l}}^{\circ}+(1-\zeta_{\mathrm{l}})\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=N_{\mathrm{ns}}/(V\varrho)$, where $\zeta_{\mathrm{l}}=1-\zeta_{\mathrm{g}}$ is the volume fraction of the dense phase (total five equations with five unknowns $\{\phi_{i\alpha}^{\mathrm{L}}\}$ and $\zeta_{\mathrm{l}}$), or by flux matching during droplet growth (Eq. (3) of the main text, for a total of five equations with five unknowns $\{\phi_{i\alpha}^{\mathrm{L}}\}$ and $\mathrm{d}R/\mathrm{d}t$).
@@
Note that Eq. (S26b) implies that $\phi_{\mathrm{ns}}(x)$ spatially rearranges in accord with the local $\phi_{\mathrm{s}}(x)$, so the positional dependence of the former is strictly acquired from the latter, $\phi_{\mathrm{ns}}(x)\to\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x))$. By multiplying each of Eqs. (S26) by $\mathrm{d}\phi_{i}(x)/\mathrm{d}x$, summing the two equations, and integrating the result from -$\infty$ to $x$, we obtain
$f(\phi_{\mathrm{s}}(x),\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x)))-\frac{\kappa_{\mathrm{s}}}{2}\left[\frac{\mathrm{d}\phi_{\mathrm{s}}(x)}{\mathrm{d}x}\right]^{2}-f(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=\frac{\mu_{\mathrm{s}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{s}}(x)-\phi_{\mathrm{s}\mathrm{g}}^{\circ}]-\frac{\mu_{\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ},\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{k_{\mathrm{B}}T}[\phi_{\mathrm{ns}}(\phi_{\mathrm{s}}(x))-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}].$ (S27)
Via a similar procedure to Sec. SI.2, the surface tension can be found as
@@
Thus, per the Gibbs phase rule, only one unknown $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$(and its leading-order finite-droplet correction, $\lambda_{\mathrm{ns}\mathrm{g}}$) remain to be determined. For brevity, since we will find $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\ll 1$, we set $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\ll e^{-\chi}$ to be simply $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}=0$ in the main text.
Lastly, we confirm that the FH surface tension (Eq. (S14)) does not change due to “labeling” inert (nonsticky) species. First, Eq. (S26b) simplifies for the FH free energy as $\phi_{\mathrm{ns}}(x)/[1-\phi_{\mathrm{s}}(x)-\phi_{\mathrm{ns}}(x)]=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}/(1-\phi_{\mathrm{s}\mathrm{g}}^{\circ}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ})$, which is the position-dependent analogue of the inert proteome/nonsticky protein equivalence argument, Eq. (S31). Inserting it in Eq. (S28) with the FH free energy (Eq. (S29)) results in Eq. (S14) with the binary-fluid’s FH free energy (Eq. (S15)). Thus, indeed the surface tension is ignorant of inert-protein labels, so $\sigma$ is still given by Eq. (S19).
@@
In this section, we detail the full derivation of the droplet growth laws in the presence of reactions. Particularly, we list our approximations, the phenomena that should be avoided in the liquid phase to enable growth, derive the flux-matching equation (Eq. (3) of the main text), extract $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ from it in the different limits of droplet size $R$, and from it the radius growth laws. Throughout, we reference the results of the FH free energy to support general intuition.
@@
Here, we construct the reaction-diffusion equations that our analysis is based on. We denote with vectors and matrices
@@
First, following linear-irreversible thermodynamics [29], the volume-fraction fields $\bar{\phi}(\mathbf{r},t)$ respond linearly to chemical potential gradients $\delta F/\delta[\varrho\bar{\phi}(\mathbf{r},t)]$ via a (normalized) mobility matrix denoted $\varrho k_{\mathrm{B}}T\bar{\bar{\Gamma}}(\bar{\phi}(\mathbf{r},t))$. Thus, the diffusive part of the dynamics is given by
@@
where $\kappa_{ij}=\delta_{i\mathrm{s}}\delta_{j\mathrm{s}}\kappa_{\mathrm{s}}$.
@@
Upon combining both reactions and diffusion, we find
@@
In the main text, we assume a one-to-one stoichiometric ratio between $\mathrm{s}$ and $\mathrm{ns}$ for the interconversion reactions. Here, we proceed with a general ratio. Since $\mathrm{s}$ is created from $\mathrm{ns}$ and vice versa, $\upsilon_{\mathrm{s}}$ and $\upsilon_{\mathrm{ns}}$ must be of opposite signs. Without loss of generality (up to rescaling of $s(\bar{\phi})$), we choose $\upsilon_{\mathrm{s}}<0$ and $\upsilon_{\mathrm{ns}}=1$; $|\upsilon_{\mathrm{s}}|$ thus determines the stoichiometric ratio. (In the main text, $\upsilon_{\mathrm{s}}=-1$.) For later convenience, we also define $\bar{\upsilon}_{\perp}$ as the orthogonal vector to $\bar{\upsilon}$, so $\bar{\upsilon}_{\perp}^{\mathrm{T}}\bar{\upsilon}=0$. Again without loss of generality, we choose it as $\bar{\upsilon}_{\perp}=(1,|\upsilon_{\mathrm{s}}|)^{\mathrm{T}}$ (whereas $\bar{\upsilon}=(-|\upsilon_{\mathrm{s}}|,1)^{\mathrm{T}}$).
@@
To make analytical progress from Eq. (S38), we introduce a number of standard simplifications [1], valid at a late stage of coarsening. First, the droplets are assumed to be sparse (the interdroplet distances are much greater than their radii), so we consider a single phase-separated droplet as if it is in isolation, diffusively exchanging material with the slightly-supersaturated dilute phase, $0<\phi_{\mathrm{s}\mathrm{g}}^{\circ}\lesssim\varphi^{\infty}_{\mathrm{s}}\ll 1$. Assuming a large surface tension, the droplets are spherical at all times, with radius $R(t)$ much larger than the interface width $\xi$ (defined in Sec. SI.3). We split the volume fraction fields to droplet interior (liquid) and exterior (gas), $\bar{\phi}_{\mathrm{l}}(r,t)=\bar{\phi}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|<R(t),t)$ and $\bar{\phi}_{\mathrm{g}}(r,t)=\bar{\phi}(|\mathbf{r}-\mathbf{r}_{\mathrm{CM}}|>R(t),t)$, respectively, where $\mathbf{r}_{\mathrm{CM}}$ is the droplet-center coordinate. Assuming the reactions are slow enough not to interfere with surface thermodynamics [31, 30], the two volume fractions are always equilibrated at the interface, $\bar{\phi}_{\alpha}(R(t),t)=\bar{\phi}_{\alpha}^{\mathrm{L}}(R(t))$ ($\alpha=\mathrm{l},\mathrm{g}$). Lastly, the droplets grow quasistatically, whereby the volume fraction fields $\{\bar{\phi}_{\alpha}(r;R)\}$ are found for a quasi-steady state of a fixed radius $R$, and then droplet growth rate is inferred from the resulting steady diffusive flux, $\bar{J}_{\alpha}(R)$.
SII.2 Droplet growth in terms of particle flux
Prior to continuing, we derive Eq. (3) of the main text, relating the droplet-radius growth rate to the particle fluxes. We integrate Eq. (S38) (assuming spherical symmetry) over the entire container,
where in the last equality we recalled the fast-surface-thermodynamics boundary condition, $\bar{\phi}_{\alpha}(R,t)=\bar{\phi}_{\alpha}^{\mathrm{L}}(R)$ ($\alpha=\mathrm{l},\mathrm{g}$), and $\bar{J}_{\alpha}(R)$ are the outward steady fluxes on either side of the interface of a droplet whose radius is $R$. Upon comparing Eqs. (S39) and (S40), we cancel out the reaction term $\bar{\upsilon}\int_{0}^{\infty}\mathrm{d}rr^{2}s(\bar{\phi}(r,t))$ and isolate $\mathrm{d}R(t)/\mathrm{d}t$
$\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\bar{J}_{\mathrm{l}}(R(t))-\bar{J}_{\mathrm{g}}(R(t)).$ (S41)
This is Eq. (3) of the main text. To be able to compute $\mathrm{d}R/\mathrm{d}t$, we must evaluate the steady interface fluxes from the field-theory equations.
SII.3 Interim simplifications
In this work, we consider a regime where the deviations of the volume fraction fields from their thermodynamic-equilibrium values are assumed to be small, so linearizing Eq. (S38) deep in each bulk might be justified. This strategy is often employed in the literature [62, 3], as it allows one to find simple explicit expressions for the fluxes. While this strategy might be justified for the dilute phase, we encounter various destabilizing phenomena in the dense phase rendering the linearization inaccurate. Thus, for completeness, we now show how linearization is performed in a general case. We then perform it explicitly for the dilute phase in Sec. SII.4 for which it is valid; with it, we find the dilute-phase fluxes. Afterwards, in Sec. SII.5 we perform the linearization in the dense phase, identify the problematic dense-phase phenomena, and then present the approach we use in practice to compute the dense-phase flux without linearization in Sec. SII.5.4.
@@
For brevity, we will omit the reminder of the parametric dependence on $R$ henceforth. Note that, in accord with Sec. SI.4, the volume fractions $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}$ ($\simeq 1$), $\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}$ ($\ll 1$), and $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}$ ($\to 0$) are known explicitly, whereas $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is still unknown. As mentioned there and in the main text, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ will be determined by equating the two equations for $\mathrm{d}R/\mathrm{d}t$, obtained from the sticky and nonsticky fluxes (Eq. (S41) or Eq. (3) of the main text; this is done in Sec. SII.6 below). To obtain the equation for $\mathrm{d}R/\mathrm{d}t$, we regard $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ as a known parameter, and express the flux $J_{\mathrm{ns}}$ in term of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. The linearization of Eq. (S42) is thus an ansatz for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ — that $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\simeq\varphi_{\mathrm{ns}}^{\infty}$ — which we make to be able to compute the flux $J_{\mathrm{ns}}$ in terms of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. Once all calculations are complete and $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is found, we confirm that indeed the resulting $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is close to $\varphi_{\mathrm{ns}}^{\infty}$.